OpenStudy (anonymous):

find the point where the normal line to y = x + root x at (4,6) crosses the y-axis.

OpenStudy (phi):

did you find the equation for the slope of the curve?

OpenStudy (anonymous):

wait.. i'll solve for the slope

OpenStudy (anonymous):

take the derivative replace \(x\) by \(4\) that gives you the slope take the negative reciprocal, that give you the slope of the normal line use the point slope formula with the point \((4,6)\) to find the equation of the normal line \(y=mx+b\) then \(b\) is where it crosses the y axis

OpenStudy (anonymous):

the value of b is the answer then?

OpenStudy (anonymous):

yes, that is the y intecept

OpenStudy (anonymous):

ok.. thanks.. i'll try solving for it

OpenStudy (phi):

Here is a plot of the scenario

OpenStudy (anonymous):

thanks! what i got was 31/5 .. is this correct?

OpenStudy (phi):

how ?

OpenStudy (anonymous):

i follow the steps satelite gave

OpenStudy (phi):

I think you wandered off the path...

OpenStudy (anonymous):

why??

OpenStudy (phi):

see the plot I posted up above

OpenStudy (anonymous):

where is the normal line there?

OpenStudy (phi):

the red line

OpenStudy (anonymous):

so what do you think is wrong with my answer? is the steps given by satelite somehow have mistakes? i just followed it

OpenStudy (phi):

what did you get for the slope ?

OpenStudy (phi):

first, what did you get for the derivative?

OpenStudy (anonymous):

owh wait.. i got a mistake

OpenStudy (anonymous):

b = 46/5 .. i believe that is the correct answer now?

OpenStudy (phi):

yes, 9.2 which matches the plot

OpenStudy (anonymous):

9.3 :)) thanks so much! help me again.. i'll post a question :)

OpenStudy (phi):

9.3 ?

OpenStudy (anonymous):

yes.. divide 46 by 5 right ? then it will be 9... remainder is 15.. that will be 3.. so 9.3?

OpenStudy (phi):

5*9= 45

OpenStudy (anonymous):

lol.. i got a mistake again. sorry.. yes.. 9.2